OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -3 and 3x 12 Answer A number line with an open circle on -1, a closed circle on 4, and shading in between. A number line with a closed circle on -1, shading to the left, and an open circle on 4, shading to the right. A number line with a closed circle on -1, an open circle on 4, and shading in between. A number line with an open circle on -1, shading to the left, and a closed circle on 4, shading to the right.

jimthompson5910 (jim_thompson5910):

What's in between the 3x and the 12?

OpenStudy (anonymous):

damnit i always forget... theres a <=

jimthompson5910 (jim_thompson5910):

x - 2 > -3 and 3x <= 12 x > -3+2 and x <= 12/3 x > -1 and x <= 4 -1 < x and x <= 4 -1 < x <= 4 The graph of the solution set on the number line is attached. Looking at the attached graph, we see that there's an open circle at -1 and an closed circle at 4. Everything in between is shaded. So the answer is choice A